Zach has sold 400 shares of IBM at $150 per share. The initial margin requirement for the short sale was 60%. The next week the share price became $160. What is Zach's percentage margin at this point?

50%

43.8%

62.5%

40%
Answer #1

Answer: Correct answer is 50%
Margin call price=Price*(1+ initial margin)/(1+maintenance margin)
We need to determine the maintenance margin.
Price=$150
Initial margin=60%
Suppose the margin call price=$160
Substituting the values in the equation, we get
160=150*(1+60%)/(1+maintenance margin)
=>(1+maintenance margin)=150*(1+60%)/160
=>Maintenance margin=150*(1+60%)/160-1=240/160-1=0.5 or 50%

So, Zach's percentage margin at the given point=50%.